OpenStudy (anonymous):

Two functions are shown in the table below. Function 1 2 3 4 5 6 f(x) = −x2 + 4x + 12 g(x) = x + 8 Complete the table on your own paper, then select the value that is a solution to f(x) = g(x). x = 3 x = 4 x = 5 x = 6

OpenStudy (anonymous):

I'll show you how to put in a value for each function, then you can fill in the table. For 1, plug in 1 for x in f(x) and g(x) f(1) = -(1)² + 4(1) + 12 = 15 g(1) = 1 + 8 = 9 then put them in the table